# Reinforcement Learning with Generalizable Gaussian Splatting

### v1

@@ VI Conclusion
In this work, we propose to use a novel environmental representation i.e. 3DGS, generally for RL algorithms. Traditional 3DGS is not suited to be an environment representation due to the need for per-scene optimization. To address this, we propose a learning framework that directly predicts 3D Gaussians from visual observations in a generalizable manner, which contains three main procedures, a depth estimator, a Gaussian regressor, and a Gaussian refinement module. This framework can effectively inject trained 3DGS priors into RL learning paradigms. To evaluate the effectiveness of the novel scene representation, we compare it with other explicit vision representations on the Robomimic platform with four different tasks and across three different RL algorithms. The results show that our generalizable GS representation overall outperforms all counterparts, and improves the Success Rate by 10, 44, and 15 respectively on the most challenging task.

### v3

@@ VI Conclusion
We propose to use a novel scene representation i.e. 3DGS, for vision-based RL algorithms. Traditional 3DGS is not suited to be an environment representation due to the need for per-scene optimization. To address this, we introduce the learning framework that directly predicts 3D Gaussians from visual observations in a generalizable manner, which contains three main procedures, a depth estimator, a Gaussian regressor, and a Gaussian refinement module. This framework can effectively capture task-specific priors of image-to-local structures. To evaluate the effectiveness of the novel scene representation, we compare it with other explicit vision representations on the Robomimic platform with four different tasks and across three different RL algorithms. The results show that our generalizable GS representation overall outperforms these counterparts, and improves the Success Rate by 10, 44, and 15 respectively on the most challenging task.